# PO-Flow: Flow-based Generative Models for Sampling Potential Outcomes and Counterfactuals

### v1

@@ B.1 Synthetic Causal Datasets
Moreover, under the following dataset synthesis assumptions: (1) fully specified outcome model $Y^{(A)}=f(X,A)+\epsilon$; (2) shared noise across treatment arms for each individual; and (3) invertibility of the structural function such that, given $X$, $A$, and $Y^{(A)}$, one can solve for $\epsilon$ — we can synthesize the counterfactual outcome through the following procedure:
@@
2. Action: intervene by changing the treatment assignment to $1-A$;
@@
B.2 Counterfactual Image Experiments
To intuitively visualize PO-Flow’s counterfactual performance and evaluate its effectiveness on high-dimensional data, we train PO-Flow on the latent space of a pre-trained Variational Autoencoder (VAE) [36] using the CelebA dataset.
The CelebFaces Attributes Dataset (CelebA) comprises 202,599 face images, each with a resolution of $178\times 218$ pixels. Each image is annotated with 40 binary facial attributes. In our experiments, we selected the male/female attribute as the treatment variable $A\in\{\text{male},\text{female}\}$, and defined covariates $X\in\mathbb{R}^{6}$ to represent features such as hair, eyebrows, bangs, and lips.
The VAE is pre-trained on the full CelebA dataset with a 512-dimensional latent space. Each $178\times 218$ face image is then encoded into this latent space, yielding the training target $Y\in\mathbb{R}^{512}$. PO-Flow is subsequently trained on the triplet $(Y,X,A)$ to generate counterfactual male face representations given original female faces.
B.3 Neural Network Structure and Implementation Information
The neural network used in experiments on causal datasets (except the CelebA image task) takes as input $y(t)$, $x$, $t$, and $a$. The model is composed of the following components: (1) an outcome embedding layer that projects $y(t)$ to the same dimension as the concatenated $(x,a)$; (2) a FiLM layer that applies feature-wise affine transformations to the embedded $y(t)$, conditioned on $(x,a)$; (3) two residual blocks, each consisting of a gated two-layer MLP conditioned on $(x,a)$, with residual averaging and skip connections; and (4) a projection layer that maps the final hidden representation to two velocity vectors $v_{0}$ and $v_{1}$, corresponding to treatments 0 and 1.
The neural network used for the counterfactual image generation task has an architecture composed of the following components: (1) a conditioning MLP that encodes the pair $(t,a)$ into a hidden feature vector via two linear layers and a ReLU activation; (2) an input layer that takes the concatenation of $y(t)$ and $(t,a)$, followed by a ReLU-activated linear transformation; (3) four residual blocks, each composed of two linear layers with a ReLU nonlinearity, where the conditioning feature is injected additively at each block; and (4) an output linear layer that maps the final hidden representation back to the latent outcome dimension.
All experiments are conducted on a V100 GPU with 16 GB of memory. The training time for causal datasets (excluding CelebA counterfactual face image generation) depends on the dataset and batch sizes. For example, training on the ACIC 2018 dataset with a batch size of 256 typically takes 2–4 seconds per epoch. Log-density evaluation, as discussed in Section 4.3, involves divergence integration and requires significantly more time than other test metrics.
B.4 Runge-Kutta numerical integration method
We apply the classical 4th-order Runge-Kutta method to solve the Neural ODE in the forward process of PO-Flow. Given a factual outcome $y^{(a)}$, we aim to compute the latent embedding $z_{y}=y(1)$ by integrating the velocity field $v_{\theta}$ from $t=0$ to $t=1$, with initial condition $y(0)=y^{(a)}$. Let $h$ denote the step size, and define discrete time steps $t_{k}=kh$, for $k=0,1,\dots,T$, such that $Th=1$. The update rule at each step is:
@@ B.4 Runge-Kutta numerical integration method
with $y(0)=y^{(a)}$ and $y(1)=z_{y}$. This produces a numerically integrated trajectory from the observed sample to its latent representation.
B.5 Hutchinson trace estimator
@@
where $p(\epsilon)$ is a distribution in $\mathbb{R}^{d_{Y}}$ satisfying $\mathbb{E}[\epsilon]=0$ and Cov$(\epsilon)=I$ (e.g., a standard Gaussian). This approximation becomes exact as $\sigma\to 0$.
B.6 Adapted Implementation Details for DiffPO and GANITE Counterfactuals
@@
  4:       $z_{t+1}\leftarrow\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}z_{t}+f_{\theta}(z_{t},t|x,a)\left(\sqrt{1-\alpha_{t+1}}-\sqrt{\frac{\alpha_{t+1}(1-\alpha_{t})}{\alpha_{t}}}\right)$
  5:     $z_{y}\leftarrow z_{T}$
@@
Appendix C Additional Experimental Results
C.1 RMSE Convergence
@@
PO-Flow demonstrates rapid convergence across datasets, requiring only around 250 training iterations (not epochs) with a batch size of 200 per iteration. This highlights the model’s strong training efficiency.
C.2 Additional counterfactual images
Figure 6: PO-Flow counterfactual image generation from female to male faces. $(a)$: Each row shows a factual outcome $y^{(a)}$ (1st column), its encoded latent $z_{y}$ (2nd column), and reverse-time decoding at $t=2/3$, $t=1/3$, and $t=0$ (3rd–5th columns), yielding the counterfactual $\hat{y}^{(1-a)}$. $(b)$: Counterfactuals decoded from perturbed latents $z_{y}+\delta$, with $\delta\sim\mathcal{N}(0,0.2I)$.

### v2

@@ B.1 Synthetic Causal Datasets
Moreover, under the following dataset synthesis assumptions: (1) fully specified outcome model $Y^{(A)}=f(X,A)+\epsilon$; (2) shared noise across treatment arms for each individual; and (3) invertibility of the structural function such that, given $X$, AA, and $Y^{(A)}$, one can solve for $\epsilon$ — we can synthesize the counterfactual outcome through the following procedure:
@@
2. Action: intervene by changing the treatment assignment to 1A1-A;
@@
B.2 Neural Network Structure and Implementation Information
The neural network used in experiments on causal datasets takes as input $y(t)$, $x$, $t$, and $a$. The model is composed of the following components: (1) an outcome embedding layer that projects $y(t)$ to the same dimension as the concatenated $(x,a)$; (2) a FiLM layer that applies feature-wise affine transformations to the embedded $y(t)$, conditioned on $(x,a)$; (3) two residual blocks, each consisting of a gated two-layer MLP conditioned on $(x,a)$, with residual averaging and skip connections; and (4) a projection layer that maps the final hidden representation to two velocity vectors $v_{0}$ and $v_{1}$, corresponding to treatments 0 and 1.
All experiments are conducted on an A100 GPU with 16 GB of memory. The training time for causal datasets depends on the dataset and batch sizes. For example, training on the ACIC 2018 dataset with a batch size of 256 typically takes 2–4 seconds per epoch. Log-density evaluation, as discussed in Section 4.3, involves divergence integration and requires significantly more time than other test metrics.
B.3 Runge-Kutta numerical integration method
We apply the classical 4th-order Runge-Kutta method to solve the Neural ODE in the forward process of PO-Flow. Given a factual outcome $y^{(a)}$, we aim to compute the latent embedding $z=y(1)$ by integrating the velocity field $v_{\theta}$ from $t=0$ to $t=1$, with initial condition $y(0)=y^{(a)}$. Let $h$ denote the step size, and define discrete time steps $t_{k}=kh$, for $k=0,1,\dots,T$, such that $Th=1$. The update rule at each step is:
@@ B.3 Runge-Kutta numerical integration method
with $y(0)=y^{(a)}$ and $y(1)=z$. This produces a numerically integrated trajectory from the observed sample to its latent representation.
B.4 Hutchinson trace estimator
@@
[…]
B.5 Adapted Implementation Details for DiffPO and GANITE Counterfactuals
@@
 4:     $z_{t+1}\leftarrow\sqrt{\frac{\alpha_{t+1}}{\alpha}}z+f_{\theta}(z,t|x,a)\left(\sqrt{1-\alpha_{t+1}}-\sqrt{\frac{\alpha_{t+1}(1-\alpha)}{\alpha}}\right)$
 5:    $z\leftarrow z_{T}$
@@
As shown in Figure 2, standard diffusion models use noisy and stochastic forward processes that may degrade the information in $z$, producing reverse samples that resemble interventional potential outcomes rather than counterfactuals.
In practice, probability flow ODEs [53, 11] or DDIM [52] can be applied to obtain a velocity field equivalent to the ODE derived from the trained diffusion score/noise, under which diffusion and flow matching are theoretically equivalent.
Appendix C Comparisons to BGM settings
Bijective Generation Mechanisms (BGM). The BGM framework [43] shows that if the true structural mechanism $f$ is in the BGM class (bijective/strictly monotone in the exogenous noise), then any learned mechanism $\hat{f}$ that (i) is also in the BGM class and (ii) matches the observed distribution—i.e., for the same parents value $X$ and action $A$,
$\hat{f}_{X}(A,U)\ \stackrel{{\scriptstyle d}}{{=}}\ f_{X}(A,U)\quad(\text{equivalently }(\hat{f}(X,\cdot))_{\#}P_{U}=(f(X,\cdot))_{\#}P_{U}),$
yields the same counterfactuals as $f$. Their result is model-agnostic, at the class level.
PO-Flow. Our findings are specific to a particular methodology and architecture, as this is a methodology paper where we analyze continuous normalizing flows (CNFs). With the additional assumption (A3) specific to PO-Flow, Proposition 4.3 shows that the PO-Flow encoded latent $Z$ is bijective in the exogenous noise $U$. Consequently, Corollary 4.5 proves that the encode–decode procedure recovers the true counterfactual.
Relationships. Because of Proposition 4.3 (enabled by the CNF), PO-Flow implements a bijective-in-noise mechanism and is thus comparable to BGM in the sense that $\hat{f}$ is bijective in $U$. However, BGM additionally requires observational distribution matching to obtain class-level identifiability, whereas the proofs of Proposition 4.3 and Corollary 4.5 do not assume such matching.
Alternative route. The counterfactual recovery result (Corollary 4.5) can alternatively be established by imposing an additional assumption:
(A4) Observational matching. For each $(x,a)$, the PO-Flow–induced observational law matches the true one, i.e.,
$(\hat{f}_{X}(a,\cdot))_{\#}P_{U}\;=\;(f_{X}(a,\cdot))_{\#}P_{U}.$
Under (A1)–(A3), Proposition 4.3 establishes a continuously differentiable bijection $\psi_{X}:U\to Z$ (independent of $A$), so the induced mechanism $\hat{f}_{X}$ implemented by PO-Flow is bijective in the exogenous noise, i.e., PO-Flow lies in the BGM class. With the additional assumption (A4), we can therefore directly invoke the counterfactual recovery result of the BGM framework [43].
Appendix D Additional Experimental Results
D.1 Empirical Validation of Assumption (A3)
We test (A3): $p_{\theta}(Z\mid X,A)=q(Z)=\mathcal{N}(0,1)$, using a two-sample Maximum Mean Discrepancy (MMD). Specifically, we test the joint factorization $p(Z,X,A)=q(Z)p(X,A)$ by computing the MMD between the empirical joint sample $\{(z_{i},x_{i},a_{i})\}$ and a synthetic joint sample $\{(z_{i}^{\prime},x_{i},a_{i})\}$, where $z_{i}^{\prime}\sim q=\mathcal{N}(0,1)$.
We employ a product kernel $k((z,x,a),(z^{\prime},x^{\prime},a^{\prime}))=k_{Z}(z,z^{\prime})\cdot k_{X}(x,x^{\prime})\cdot k_{A}(a,a^{\prime})$, where $k_{Z}(z,z^{\prime})$ and $k_{X}(x,x^{\prime})$ are radial basis function (RBF) mixtures and $k_{A}(a,a^{\prime})=\mathbf{1}[a=a^{\prime}]$ is the indicator kernel for discrete actions:
$k_{Z}(z,z^{\prime})$ $=\exp\!\left(-\frac{\|z-z^{\prime}\|^{2}}{2\sigma_{Z}^{2}}\right),$ (49)
$k_{X}(x,x^{\prime})$ $=\exp\!\left(-\frac{\|x-x^{\prime}\|^{2}}{2\sigma_{X}^{2}}\right).$
To set the bandwidths $\{\sigma_{Z}\}$ and $\{\sigma_{X}\}$, we first compute the pooled pairwise distances on the union of the observed and synthetic sets (e.g., $\tilde{Z}=Z\cup Z^{\prime}$). We then calculate the median of these distances, and define the bandwidths as $\tfrac{1}{2}\,\text{median}$. The empirical MMD is given as:
$\widehat{\mathrm{MMD}}^{2}_{u}=\frac{1}{n(n-1)}\sum_{i\neq j}k(u_{i},u_{j})+\frac{1}{n(n-1)}\sum_{i\neq j}k(v_{i},v_{j})-\frac{2}{n(n-1)}\sum_{i\neq j}k(u_{i},v_{j}),$ (50)
with $u_{i}=(z_{i},x_{i},a_{i})$, $v_{i}=(z_{i}^{\prime},x_{i},a_{i})$.
For more reliable comparisons, we additionally sample two ground-truth sets, $\{(z_{i}^{\prime},x_{i},a_{i})\}$ and $\{(z_{i}^{\prime\prime},x_{i},a_{i})\}$, where both $z_{i}^{\prime}\sim\mathcal{N}(0,1)$ and $z_{i}^{\prime\prime}\sim\mathcal{N}(0,1)$. The test is conducted on each dataset considered in the paper, and the aggregated results are summarized in the table below.
Table 7: Empirical validation of Assumption (A3), $p_{\theta}(Z\mid X,A)=q(Z)=\mathcal{N}(0,1)$. We report two-sample MMD values comparing the empirical joint distribution $\{(z_{i},x_{i},a_{i})\}$ with synthetic samples $\{(z_{i}^{\prime},x_{i},a_{i})\}$ across datasets. Lower MMD indicates closer agreement with the assumed factorization $p(Z,X,A)=q(Z)p(X,A)$.
ACIC 2016 ACIC 2018 IHDP IBM
PO-Flow $4.9\times 10^{-2}$ $7.2\times 10^{-3}$ $2.7\times 10^{-2}$ $6.8\times 10^{-2}$
True $4.1\times 10^{-2}$ $5.9\times 10^{-3}$ $2.5\times 10^{-2}$ $5.3\times 10^{-2}$
As shown in Table 7, PO-Flow achieves small two-sample joint MMD values, which are comparable to those obtained from the ground-truth independent samples $\{(z_{i}^{\prime},x_{i},a_{i})\}$ and $\{(z_{i}^{\prime\prime},x_{i},a_{i})\}$. This provides empirical support for the validity of Assumption (A3).
D.2 Identifying the Most Likely Potential Outcomes via Log-Density
As shown in Proposition 4.7 and Section 5.3.1, we can naturally select the most likely potential outcome based on log-density. Table 8 reports the RMSE of PO-Flow when using the most likely PO selected by the algorithm. Specifically, for each sample we simulate 100 POs with PO-Flow, select the one with the highest log-density, and compute the results accordingly.
Table 8: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets. For PO-Flow, results are based on the most likely PO selected from 100 simulated outcomes using log-density. All values are averaged over 10-fold cross-validation.
ACIC 2018 IHDP IBM
PO-Flow $\mathbf{0.42_{\pm.07}}$ $\mathbf{0.54_{\pm.10}}$ $\mathbf{0.96_{\pm.09}}$ $\mathbf{1.05_{\pm.10}}$ $\mathbf{0.84_{\pm.06}}$ $\mathbf{0.90_{\pm.08}}$
INFs $0.69_{\pm.16}$ $0.78_{\pm.17}$ $1.02_{\pm.10}$ $1.20_{\pm.12}$ $1.52_{\pm.24}$ $1.57_{\pm.28}$
S-learner $0.54_{\pm.09}$ $0.57_{\pm.11}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $01_{\pm.17} 1.1$$16_{\pm.19}$
CEVAE $0.83_{\pm.17}$ $0.85_{\pm.17}$ $1.18_{\pm.15}$ $1.36_{\pm.18}$ $0.96_{\pm.12}$ $0.98_{\pm.12}$
CFR $0.94_{\pm.10}$ $0.98_{\pm.12}$ $10_{\pm{.20}} 1.1$$18_{\pm{.20}}$ $2.09_{\pm.45}$ $2.17_{\pm.48}$
Compared with the averaged RMSE over all POs generated by PO-Flow in Table 1, selecting the most likely POs based on log-density yields smaller RMSEs, providing more confident and closer estimates.
D.3 Convergence and Computational Costs
(a) Convergence of MSE and RMSE for predicted potential outcomes over the training iterations on the ACIC 2018, IHDP, and IBM datasets.
@@ D.3 Convergence and Computational Costs
Figure 5(a) shows the convergence of RMSE on multiple datasets. PO-Flow demonstrates rapid convergence across datasets, requiring only around 250 training iterations (not epochs) with a batch size of 200 per iteration. This highlights the model’s strong training efficiency.
Table 9: Comparison of model size, training time untill convergence, and sampling cost (for 500 samples).
Methods # Params Training Time Sampling
PO-Flow $32,394$ $\mathbf{0.96}$min $\mathbf{0.58}$s
DiffPO $176,372$ 5.755.75min $4.98$s
INFs $138,172$ 5.095.09min $0.74$s
GANITE $167,381$ 4.934.93min $0.80$s
CEVAE $97,864$ 4.644.64min $0.47$s
In addition, Table 9 compares model size, training time to convergence, and sampling time (for 500 samples) across PO-Flow and other deep generative baselines. All experiments were conducted on a single A100 GPU. Notably, PO-Flow attains superior results with the fewest neural network parameters, leading to faster training and substantially reduced sampling time compared with diffusion-, discrete flow-, and GAN-based methods.